The nth term t(n),of a sequence is giving by t(n)=p+nq,where p and q are constants.find the sum of the first 6 terms in terms of p and q.
Answer by Edwin McCravy(20056)

The nth term t(n),of a sequence is giving by t(n)=p+nq,where p and q are
constants.find the sum of the first 6 terms in terms of p and q.
t(n)=p+nq


t(1) =  p + 1q
t(2) =  p + 2q
t(3) =  p + 3q
t(4) =  p + 4q
t(5) =  p + 5q
t(6) =  p + 6q
--------------
 sum = 6p + 21q

Edwin